# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

### v1

@@ 5 Tasks Addressed by Multimodal RAGs
Multimodal RAG systems extend RAG beyond unimodal settings to tasks requiring cross-modal integration, enhancing performance across modalities such as text, images, and audio. In content generation, these models enhance image captioning Zhi Lim et al. (2024); Hu et al. (2023); Rao et al. (2024) and text-to-image synthesis Yasunaga et al. (2023); Chen et al. (2022b) by retrieving relevant contextual information. They also improve coherence in visual storytelling and ensure factual alignment in multimodal summarization Tonmoy et al. (2024). In knowledge-intensive applications, multimodal RAG supports open-domain and knowledge-seeking question answering Chen et al. (2024d); Ding et al. (2024b); Yuan et al. (2023), video-based QA Luo et al. (2024b), and fact verification Khaliq et al. (2024), grounding responses in retrieved knowledge and thereby mitigating hallucinations. Cross-modal retrieval further advances zero-shot image–text retrieval Yang et al. (2024); Dong et al. (2024b). Additionally, the recent incorporation of chain-of-thought reasoning Zhai (2024); Khaliq et al. (2024) has enhanced its ability to support complex problem solving and logical inference. Finally, integrating multimodal RAG into interactive agents and AI assistants such as Gemini Team et al. (2024) enables natural language-driven visual search, document understanding, and multimodal reasoning. The taxonomy of application domains can be seen in Figure 3. The following sections explore domain-specific adaptations of these techniques in greater depth.
@@ 6 Open Problems and Future Directions
Generalization, Explainability, and Robustness
@@
Multimodal RAG systems often struggle with domain-specific training and exhibit modality biases, frequently over-relying on text for both retrieval and generation Winterbottom et al. (2020). Explainability remains a key challenge, as these systems typically fail to attribute answers to precise sources. Current methods often cite entire documents or large visual regions as source attribution rather than identifying the exact part of an image, speech, or other modality that led to the answer Ma et al. (2024b); Hu et al. (2023). In addition, the interplay between modalities affects the quality of outcomes produced by these models; for example, answers derived solely from text sources may differ in quality compared to those requiring a combination of text and image inputs Baltrusaitis et al. (2019). They are also vulnerable to adversarial perturbations, such as misleading images influencing textual outputs, and their performance can degrade when relying on low-quality or outdated sources Chen et al. (2022b). While the trustworthiness of unimodal RAGs has been studied Zhou et al. (2024d), enhancing the robustness of multimodal RAGs remains an open challenge and a promising research direction.
Reasoning, Alignment, and Retrieval Enhancement
@@
Multimodal RAGs struggle with compositional reasoning, where information from different modalities must be logically integrated to generate coherent and contextually rich outputs. While cross-modal techniques such as Multimodal-CoT have been proposed Zhang et al. (2023), further innovations are needed to improve the coherence and contextual relevance of multimodal outputs. Enhancing modality alignment and retrieval strategies, particularly for entity-aware retrieval, is essential. Moreover, despite the potential of knowledge graphs to enrich cross-modal reasoning, they remain largely underexplored in multimodal RAGs compared to text-based RAGs Zhang et al. (2024g); Procko and Ochoa (2024). Retrieval biases such as position sensitivity Hu et al. (2024c), redundant retrieval Nan et al. (2024), and biases propagated from training data or retrieved content Zhai (2024), pose significant challenges that require further attention. Another promising direction is developing a unified embedding space for all modalities, enabling direct multimodal search without intermediary conversion models (e.g., ASRs). Despite some progress, mapping multimodal knowledge into a unified space remains an open challenge with significant potential.
Agent-Based and Self-Guided Systems
@@
Recent trends indicate a shift towards agent-based multimodal RAGs that integrate retrieval, reasoning, and generation across diverse domains. Unlike static RAG systems, future multimodal RAGs should integrate interactive feedback and self-guided decision-making to iteratively refine outputs. Furthermore, existing feedback mechanisms often fail to accurately identify whether issues with responses arise during retrieval, generation, or other stages Dong et al. (2024b). The incorporation of reinforcement learning and end-to-end human-aligned feedback into multimodal RAGs remains largely unexplored but holds significant potential for enhancing these systems. These methods could enable multimodal RAGs to assess whether retrieval is necessary, evaluate the relevance of retrieved content, and dynamically determine the most suitable modalities for response generation. Achieving robust support for any-to-any modality is essential for adaptability in open-ended tasks Wu et al. (2024b). Future multimodal RAGs should incorporate data from diverse real-world sources, such as environmental sensors, alongside traditional modalities like text and images, to enhance situational awareness. This progression aligns with the trend toward embodied AI, where models integrate knowledge with physical interaction, enabling applications in robotics, navigation, and physics-informed reasoning. Bridging retrieval-based reasoning with real-world agency brings these systems closer to AGI.
Long-Context Processing, Scalability, and Personalization
@@
High computational costs in video frame sampling and memory bottlenecks in processing multi-page documents with images remain key challenges in long-context processing. Fixed extraction rates often fail to capture relevant video frames, necessitating dynamic adjustment based on content complexity and movement Kandhare and Gisselbrecht (2024). Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers highlight the need for more scalable architectures. Personalization mechanisms, such as adapting retrieval to user-specific contexts like medical history, remain underexplored. As personalization mechanisms evolve, ensuring privacy and mitigating the risks of sensitive data leakage in multimodal outputs remain critical challenges. Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples hinders robust evaluation.
@@ 7 Conclusion
In this study, we take a deep dive into the realm of multimodal Retrieval-Augmented Generation (RAG) literature. Specifically, we explore and categorize key advancements across different aspects of multimodal RAG systems, including retrieval, multimodal fusion, augmentation, generation, and training strategies. Additionally, we examine the tasks these systems address, their domain-specific applications, and the datasets, benchmarks, and evaluation methods. We also discuss open challenges and limitations in current approaches. We hope it could encourage future research in the domain, particularly in improving cross-modal reasoning and retrieval, developing agent-based interactive systems, and advancing unified multimodal embedding spaces.
@@ 8 Limitations
Furthermore, multimodal RAG is a rapidly evolving field with many open questions, such as optimizing fusion strategies for diverse modalities and addressing scalability challenges. As new paradigms emerge, our taxonomy and conclusions will inevitably evolve. To address these gaps, we plan to continuously monitor developments and update this survey and the corresponding GitHub repository to incorporate overlooked contributions and refine our perspectives.
@@ Appendix A Taxonomy
In this section, we present a comprehensive taxonomy of multimodal retrieval-augmented generation (RAG) systems, along with a classification of their application domains. Figures 2 and 3 summarize these taxonomies, respectively.
@@
Figure 3
presents the taxonomy of application domains for multimodal RAG systems. The identified domains include healthcare and medicine, software engineering, fashion and e-commerce, entertainment and social computing, and emerging applications. This classification offers a concise overview of the diverse applications and serves as a framework for the more detailed analyses that follow.
@@ Appendix B Dataset and Benchmark
  • Category: This column categorizes each dataset or benchmark based on its primary domain or modality. The datasets are grouped into eight categories: Image–Text General, Video–Text, Audio–Text, Medical, Fashion, 3D, Knowledge & QA, and Other. The benchmarks are grouped into three categories: Vision-Focused, Cross-Modal and Reasoning-Focused, and Text and Knowledge-Focused. This classification facilitates a clearer understanding of each dataset or benchmark’s role within a multimodal framework.
@@
ScienceQA Lu et al. (2022) High–school science curriculum dataset with 12K training samples and over 21K questions across various subjects. Text ScienceQA
@@
MMBench Liu et al. (2025) Bilingual benchmark for assessing the multimodal capabilities of VLMs; over 3,000 multiple–choice questions covering 20 dimensions. Text MMBench
@@
Category Name Statistics and Description Modalities Link
Vision-Focused MRAG-Bench Hu et al. (2024c)
Evaluates visual information retrieval and integration;
tests robustness to irrelevant visual information and visual reasoning.
Images MRAG-Bench
VQAv2 Goyal et al. (2017a)
Contains 265,016 images with 1,105,904 questions; assesses the ability to answer questions about images,
integrating image and question text, visual reasoning, and common-sense reasoning.
Images + Text VQAv2
OK-VQA Marino et al. (2019)
Comprises 14,031 questions; tests retrieval and integration of external information,
combining image, question, and external knowledge, focusing on reasoning with external knowledge.
Images + Text OK-VQA
Cross-Modal and Reasoning-Focused M2RAG Ma et al. (2024c)
Establishes a comprehensive benchmark for multimodal retrieval-augmented generation;
integrates fine-grained text-modal and multimodal metrics to evaluate foundation models on dynamic retrieval,
multi-hop reasoning, and multimodal integration.
Images + Text M2RAG
Dyn-VQA Li et al. (2024b)
Emphasizes dynamic retrieval with rapidly changing information;
requires integration of dynamic knowledge retrieval and multi-hop reasoning, testing robustness to changing information.
Images + Text Dyn-VQA
MMBench Liu et al. (2025)
Covers various tasks, including visual question answering, image captioning, and cross-modal retrieval;
focuses on integrating information across multiple modalities (vision, text, audio) for general multimodal understanding.
Images + Text + Audio MMBench
ScienceQA Saikh et al. (2022)
Contains 21,208 questions; tests scientific reasoning abilities using multimodal contexts (text, diagrams, images);
requires deep understanding and integration of information, scientific reasoning, and diagram understanding.
Images + Diagrams + Text ScienceQA
SK-VQA Su et al. (2024c)
Offers over 2 million question-answer pairs;
focuses on synthetic knowledge generation, multimodal reasoning, and external knowledge integration;
used for training context-augmented multimodal large language models.
Images + Text SK-VQA
Text and Knowledge-Focused TriviaQA Joshi et al. (2017)
Provides 650K question-answer pairs; a reading comprehension dataset where questions are based on evidence documents;
can be adapted to multimodal RAG by providing images or other visual context related to the trivia;
focuses on knowledge retrieval and factual accuracy.
Text TriviaQA
Natural Questions Kwiatkowski et al. (2019)
Contains 307,373 training examples; questions are naturally occurring Google search queries;
adapting this for multimodal RAG involves finding or creating relevant visual contexts for the queries;
emphasizes knowledge retrieval and handling noisy or ambiguous queries.
Text Natural Questions
SMMQG Wu et al. (2024a)
Includes 1,024 question-answer pairs; focuses on synthetic data generation,
multimodal retrieval-augmented generation, and control over question style and modality.
Images + Text SMMQG

### v2

@@ 5 Tasks Addressed by Multimodal RAGs
Multimodal RAG systems extend RAG beyond unimodal settings to tasks requiring cross-modal integration, enhancing performance across modalities such as text, images, and audio. In content generation, these models enhance image captioning Zhi Lim et al. (2024); Hu et al. (2023); Rao et al. (2024) and text-to-image synthesis Yasunaga et al. (2023); Chen et al. (2022b) by retrieving relevant contextual information. They also improve coherence in visual storytelling and ensure factual alignment in multimodal summarization Tonmoy et al. (2024). In knowledge-intensive applications, multimodal RAG supports open-domain and knowledge-seeking question answering Chen et al. (2024d); Ding et al. (2024b); Yuan et al. (2023), video-based QA Luo et al. (2024b), and fact verification Khaliq et al. (2024), grounding responses in retrieved knowledge and thereby mitigating hallucinations.
Cross-modal retrieval further advances zero-shot image–text retrieval Yang et al. (2024); Dong et al. (2024b). Additionally, the recent incorporation of chain-of-thought reasoning Zhai (2024); Khaliq et al. (2024) has enhanced its ability to support complex problem solving and logical inference. Finally, integrating multimodal RAG into interactive agents and AI assistants such as Gemini Team et al. (2024) enables natural language-driven visual search, document understanding, and multimodal reasoning. The taxonomy of application domains can be seen in Figure 3. The following sections explore domain-specific adaptations of these techniques in greater depth.
@@ 6 Open Problems and Future Directions
6.1 Generalization, Explainability, and Robustness
@@
Multimodal RAG systems often struggle with domain adaptation and exhibit modality biases, frequently over-relying on text for both retrieval and generation Winterbottom et al. (2020). Explainability remains a major challenge, as these systems typically fail to attribute answers to precise sources. Current methods often cite entire documents or large visual regions as source attribution rather than identifying the exact part of an image, speech, or other modality that led to the answer Ma et al. (2024b); Hu et al. (2023).
In addition, the interplay between modalities affects the quality of outcomes produced by these models; for example, answers derived solely from text sources may differ in quality compared to those requiring a combination of text and image inputs Baltrusaitis et al. (2019). They are also vulnerable to adversarial perturbations, such as misleading images influencing textual outputs, and their performance can degrade when relying on low-quality or outdated sources Chen et al. (2022b). While the trustworthiness of unimodal RAGs has been studied Zhou et al. (2024d), enhancing the robustness of multimodal RAGs remains an open challenge and a promising research direction.
6.2 Reasoning, Alignment, and Retrieval Enhancement
@@
Multimodal RAGs struggle with compositional reasoning, where information from different modalities must be logically integrated to generate coherent and contextually rich outputs. While cross-modal techniques such as Multimodal-CoT Zhang et al. (2023b) have been proposed, further innovations are needed to improve the coherence and contextual relevance of multimodal outputs. Enhancing modality alignment and retrieval strategies, particularly for entity-aware retrieval, is essential. Moreover, despite the potential of knowledge graphs to enrich cross-modal reasoning, they remain largely underexplored in multimodal RAGs compared to text-based RAGs Zhang et al. (2024f); Procko and Ochoa (2024).
Retrieval biases such as position sensitivity Hu et al. (2024c), redundant retrieval Nan et al. (2024), and biases propagated from training data or retrieved content Zhai (2024), pose significant challenges that require further attention. Another promising direction is developing a unified embedding space for all modalities, enabling direct multimodal search without intermediary conversion models (e.g., ASRs). Despite some progress, mapping multimodal knowledge into a unified space remains an open challenge with significant potential.
6.3 Agent-Based and Self-Guided Systems
@@
Recent trends indicate a shift towards agent-based multimodal RAGs that integrate retrieval, reasoning, and generation across diverse domains. Unlike static RAG systems, future multimodal RAGs should integrate interactive feedback and self-guided decision-making to iteratively refine outputs. Existing feedback mechanisms often fail to accurately determine whether errors stem from retrieval, generation, or other stages Dong et al. (2024b). The incorporation of reinforcement learning and end-to-end human-aligned feedback into multimodal RAGs remains largely unexplored but holds significant potential for enhancing these systems. These methods could enable multimodal RAGs to assess whether retrieval is necessary, evaluate the relevance of retrieved content, and dynamically determine the most suitable modalities for response generation. Achieving robust support for any-to-any modality is essential for adaptability in open-ended tasks Wu et al. (2024b).
Future multimodal RAGs should incorporate data from diverse real-world sources, such as environmental sensors, alongside traditional modalities like text and images, to enhance situational awareness. This progression aligns with the trend toward embodied AI, where models integrate knowledge with physical interaction, enabling applications in robotics, navigation, and physics-informed reasoning. Bridging retrieval-based reasoning with real-world agency brings these systems closer to AGI.
6.4 Long-Context Processing, Efficiency, Scalability, and Personalization
@@
High computational costs in video frame sampling and memory bottlenecks in processing multi-page documents with images remain key challenges in long-context processing. Fixed extraction rates struggle to capture relevant frames, requiring adaptive selection based on content complexity and movement Kandhare and Gisselbrecht (2024). Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers emphasize the need for efficient, scalable architectures. Personalization mechanisms, such as adapting retrieval to user-specific contexts like medical history, remain underexplored. As personalization mechanisms evolve, ensuring privacy and mitigating the risks of sensitive data leakage in multimodal outputs remain critical challenges. Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples limits robust evaluation.
@@ 7 Conclusion
This study provides a comprehensive review of multimodal Retrieval-Augmented Generation (Multimodal RAG) literature. Specifically, we explore and categorize key advancements across different aspects of multimodal RAG systems, including retrieval, multimodal fusion, augmentation, generation, and training strategies. Additionally, we examine the tasks these systems address, their domain-specific applications, and the datasets, benchmarks, and evaluation methods. We also discuss open challenges and limitations in current approaches, along with promising future directions. We hope this work inspires future research, particularly in enhancing cross-modal reasoning and retrieval, developing agent-based interactive systems, and advancing unified multimodal embedding spaces.
@@ 8 Limitations
Furthermore, multimodal RAG is a rapidly evolving field with many open questions, such as optimizing fusion strategies for diverse modalities and addressing scalability challenges. As new paradigms emerge, our taxonomy and conclusions will inevitably evolve. To address these gaps, we plan to continuously monitor developments and update this survey and the corresponding repository to incorporate overlooked contributions and refine our perspectives.
9 Ethical Statement
This survey provides a comprehensive review of research on multimodal RAG systems, offering insights that we believe will be valuable to researchers in this evolving field. All the studies, datasets, and benchmarks analyzed in this work are publicly available, with only a very small number of papers requiring institutional access. Additionally, this survey does not involve personal data or user interactions, and we adhere to ethical guidelines throughout.
Since this work is purely a survey of existing literature and does not introduce new models, datasets, or experimental methodologies, it presents no potential risks. However, we acknowledge that multimodal RAG systems inherently raise ethical concerns, including bias, misinformation, privacy, and intellectual property issues. Bias can emerge from both retrieval and generation processes, potentially leading to skewed or unfair outputs. Additionally, these models may hallucinate or propagate misinformation, particularly when retrieval mechanisms fail or rely on unreliable sources. The handling of sensitive multimodal data also poses privacy risks, while content generation raises concerns about proper attribution and copyright compliance. Addressing these challenges requires careful dataset curation, bias mitigation strategies, and transparent evaluation of retrieval and generation mechanisms.
@@ Appendix A Taxonomy
In this section, we provide more details regarding the taxonomy of multimodal RAG systems, previously mentioned in Figure 2. Additionally, we present a classification of multimodal RAG application domains in Figure 3.
@@
Figure 3 presents the taxonomy of application domains for multimodal RAG systems. The identified domains include healthcare and medicine, software engineering, fashion and e-commerce, entertainment and social computing, and emerging applications. This classification offers a concise overview of the diverse applications and serves as a framework for the more detailed analyses that follow.
@@ Appendix B Dataset and Benchmark
  • Category: This column categorizes each dataset or benchmark based on its primary domain or modality. The datasets are grouped into eight categories: Image–Text General, Video–Text, Audio–Text, Medical, Fashion, 3D, Knowledge & QA, and Other. The benchmarks are grouped into two categories: Cross-Modal Understanding and Text-Focused. This classification facilitates a clearer understanding of each dataset or benchmark’s role within a multimodal framework.
@@
SpeechBrown Abootorabi and Asgari (2024) 55K paired speech-text samples; 15 categories covering diverse topics from religion to fiction. Audio, Text SpeechBrown
@@
Category
Name Statistics and Description Modalities Link
Cross-Modal Understanding MRAG-Bench Hu et al. (2024c) Evaluates visual retrieval, integration, and robustness to irrelevant visual information. Images MRAG-Bench
M2RAG Ma et al. (2024c) Benchmarks multimodal RAG; evaluates retrieval, multi-hop reasoning, and integration. Images + Text M2RAG
Dyn-VQA Li et al. (2024b) Focuses on dynamic retrieval, multi-hop reasoning, and robustness to changing information. Images + Text Dyn-VQA
MMBench Liu et al. (2025) Covers VQA, captioning, retrieval; evaluates cross-modal understanding across vision, text, and audio. Images + Text + Audio MMBench
ScienceQA Saikh et al. (2022) Contains 21,208 questions; tests scientific reasoning with text, diagrams, and images. Images + Diagrams + Text ScienceQA
SK-VQA Su et al. (2024b) Offers 2 million question-answer pairs; focuses on synthetic knowledge, multimodal reasoning, and external knowledge integration. Images + Text SK-VQA
SMMQG Wu et al. (2024a) Includes 1,024 question-answer pairs; focuses on synthetic multimodal data and controlled question generation. Images + Text SMMQG
Text-Focused TriviaQA Joshi et al. (2017) Provides 650K question-answer pairs; reading comprehension dataset, adaptable for multimodal RAG. Text TriviaQA
Natural Questions Kwiatkowski et al. (2019) Contains 307,373 training examples; real-world search queries, adaptable with visual contexts. Text Natural Questions